A person is riding in a car which is moving at a constant speed around a circular arc.




The person has an accelerometer on their mobile phone and uses it to measure a lateral (horizontal) acceleration of 5.6 m s−2.




If the car was traveling at 50 km h−1 (14 m s−1) then what was the radius of the curve that car was traveling along? (to 2 s.f)





1
Expert's answer
2022-04-11T12:04:49-0400

"a=\frac{v^2}r,\implies"

"r=\frac{v^2}a=\frac{14^2}{5.6}=35~m."